# Mixture of Multicenter Experts in Multimodal Generative AI for Advanced Radiotherapy Target Delineation

### v1

@@ Discussion
To address the challenges of variability in patient characteristics and treatment strategies across different institutions and countries, we propose a multicenter training strategy that incorporates diverse clinical approaches using Mixture of Multicenter Expert (MoME) modules within a multimodal alignment learning framework. Our center-specific adaptation aims to manage the complexity and heterogeneity of patient populations and treatment protocols. This method allows the model to learn essential knowledge from large-scale datasets for patient treatment planning while simultaneously specializing in each center’s unique patient patterns with just one-shot training across five classes, requiring only 10 additional training samples per center, including a one-shot validation dataset. The superiority of our multimodal MoME training is demonstrated not only through quantitative analyses using traditional metrics such as the Dice coefficient but also qualitatively via GPR analysis, which systematically converts center-specific practice patterns into distributions for comparison. This analysis showed that the MoME framework effectively aligns with individual center practices. Fig.3(a) further reinforces these findings by demonstrating that, when using center-specific routers, the model closely adapts to the treatment patterns of each center, and the patterns between centers with similar practices exhibit comparable trends. Beyond the ability to fine-tune using individual institutional data, our method proves highly adaptable in clinical settings where data sharing is restricted, but adaptation to new data distributions is necessary. As demonstrated in Table 4, our proposed MoME framework enables versatile zero-shot inference scenarios in clinical settings, allowing for the selection of the most relevant center-specific router by utilizing only a few sample test datasets, rather than relying on a traditional naive inference approach. This effectively addresses the long-standing challenges of deploying AI models in clinical practice by showing that the performance of the most relevant center-specific inference significantly exceeds that of the least relevant center-specific inference, while also surpassing traditional vision-only and vanilla multimodal inference with substantial gains. Furthermore, to optimize the pre-trained model for the deployed clinical settings, one- to three-shot fine-tuning using the selected center-specific router network with MoME outperforms traditional AI training mechanisms, as shown in Fig. 4. This approach proves particularly valuable for real-world applications with limited sample datasets. The experimental results, conducted with varying dataset sizes, suggest that our proposed MoME fine-tuning approach efficiently adapts to each institution’s treatment strategies with minimal multicenter data, particularly for complex tasks like radiotherapy target volume delineation, where institutional consensus is often difficult to achieve. This demonstrates the potential for AI-based models to be effectively deployed in clinical scenarios, offering a promising solution for adapting to diverse patient populations and treatment protocols in a fast and efficient manner.
Despite the advantages of our multicenter training mechanism over traditional AI methods, several limitations should be acknowledged. One primary drawback is the reliance on a limited number of samples for fine-tuning. While few-shot training demonstrated effectiveness in our experiments, the small sample size may limit the model’s robustness in more complex or diverse clinical environments. Specifically, in the closed center fine-tuning setting, our method showed relative superiority (as demonstrated in Table 2), but the overall performance remained below 0.7 on the Dice metric, indicating that the model requires further improvement before it can be used in clinical practice. Additionally, although we employed few-shot datasets from each center, especially for closed centers, future research should focus on incorporating larger and more diverse datasets to enhance the model’s robustness and generalizability. Broader validation across various cancer types and treatment plans will also be essential to confirm the wider applicability of our approach. Lastly, while we introduced the GPR metric to visualize and systematically compare the differences in practice patterns between institutions, it remains an indirect measure that captures only certain aspects of these patterns. For example, one critical factor in prostate cancer target delineation is the inclusion of the seminal vesicles. While the inclusion or exclusion of this structure may slightly affect the PTV margin, its overall impact on the GPR distribution might be minimal, meaning this aspect of practice variation may not be fully captured by the GPR analysis alone. Although visualizing the internal expert router patterns helped demonstrate MoME’s effectiveness, the GPR metric does not encompass all aspects of institutional practice patterns.
@@
In conclusion, our study marks a significant step toward enabling collaboration on multicenter datasets despite challenges associated with large-scale data collection and practical constraints across institutions. It offers an effective method for addressing variability in radiotherapy target delineation practices. Our approach significantly outperformed traditional AI training methods, demonstrating strong generalization to diverse clinical settings and adaptability to distribution shifts. This adaptability further positions the MoME framework as a promising candidate for multicenter collaborations, especially in addressing complex and often debated clinical decision-making tasks by fostering collaborative synergy and aligning with institutional unique strategies.
@@ Methods
0.6 Ethic Committee Approval.
The hospital data deliberately collected for this study were ethically approved by the Institutional Review Board of Department of Radiation Oncology at Yonsei Cancer Center, Department of Radiation Oncology at Yongin Severance Hospital, and Department of Radiation Oncology at Gangnam Severance Hospital (approval numbers of 4-2023-0179, 9-20230161, and 3-2023-0396 for each), Department of Radiation Oncology at Mayo Clinic (approval numbers of 13-005709), and Massachusetts General Hospital (MGH) (approval numbers of 2022P001512). The requirement for informed consent was waived due to the retrospective nature of the study.
0.7 Details of Target Volume for Prostate Cancer.
In radiation oncology, treatment target volumes are classified as Gross Tumor Volume (GTV), Clinical Target Volume (CTV), and Planning Target Volume (PTV). GTV represents the visible tumor, aligning with the goal of traditional segmentation to delineate observable image regions. CTV, while sometimes derived directly from GTV when a gross tumor is present, typically encompasses areas at risk for microscopic disease. Defining CTV requires consideration of multiple clinical factors, including tumor type, histological findings, cancer stage according to Tumor, Node, Metastasis classification (TNM classification) system, and, in specific cases, patient age and performance status. PTV further extends CTV to account for uncertainties in patient setup and positioning.
@@
In definitive radiotherapy, the target volume typically includes the prostate, seminal vesicles, and areas of suspected extracapsular extension [31]. For postoperative radiotherapy, the target encompasses the prostate bed and the seminal vesicle bed, with consideration of the surrounding anatomical structures based on the surgical field [32]. If lymph node involvement is confirmed, or if the patient falls within the high-risk or very high-risk groups according to NCCN guidelines [22], PNI is recommended, even in the absence of radiographic evidence of lymph node metastasis. Despite the existence of general principles guiding target volume delineation, there is considerable variability in practice, particularly in the specifics of margin sizes and the inclusion of adjacent structures in cases of suspected locoregional invasion. Intermediate-risk patients with unfavorable factors may also receive PNI depending on institutional protocols, although this is not universally applied.
To reduce variability in target volume delineation, multiple consensus guidelines have been developed for both definitive and postoperative radiotherapy in prostate cancer [31, 22]. However, significant differences between institutions persist, driven by factors such as clinician preferences, institutional protocols, educational environments, and regional practices. Complete standardization remains challenging, emphasizing the need for strategies that can adapt to the specific practices of individual institutions while maintaining high standards of care.
0.8 Details of Dataset.
@@
0.9 Implementation Details.
The detailed schematic of our proposed network is illustrated in Supplementary Fig. 1. For the image encoder/decoder for each organ extractor and PTV delineation network, we employed the 3D Residual U-Net [24]. For the large language model (LLM), we utilized the pre-trained Llama3-8B-chat [33] model. Our proposed multimodal MoME is composed of four key steps: (a) prostate segmentation, (b) interactive multimodal alignment, (c) center-specific MoME training, and (d) closed center MoME fine-tuning.
0.10 (a) Prostate Segmentation
@@
0.11 (b) Interactive Multimodal Alignment
For our multimodal AI training, we adapted a large language model (LLM) to extract clinical data. As illustrated in Supplementary Fig. 1 (b), to efficiently transfer the extensive knowledge of the LLM, we employed text prompt tuning, as previously introduced in our context-aware 3D segmentation network, LLMSeg [4]. In this approach, the frozen LLM embeds clinical data into context embeddings ${g}\in{\mathbb{R}^{1\times D}}$, where $D$ is the LLM embedding dimension. To align these context embeddings ${g}$ with the image embeddings ${f_{l}}\in\mathbb{R}^{H_{l}W_{l}S_{l}\times Ch_{l}}$, where ${f_{l}}$ is the output of the $l$-th layer of the 3D image encoder, with $H_{l}$, $W_{l}$, and $S_{l}$ representing the height, width, and slice of the image embeddings, and $Ch_{l}$ the intermediate channel dimension, we first project ${g}$ to match the dimensions of each ${f_{l}}$ using a layer-wise linear transformation. Then, these linearly projected context embeddings ${\bar{g}}_{l}\in{\mathbb{R}^{1\times Ch_{l}}}$ are subsequently processed through self-attention and cross-attention mechanisms with the image embeddings ${f_{l}}$ within two-way transformer modules of SAM [34] structure, resulting in multimodal image embeddings ${\bar{f_{l}}}\in\mathbb{R}^{H_{l}W_{l}S_{l}\times Ch_{l}}$. In contrast to the original LLMSeg, for multicenter training, we further introduced a mixture of multicenter experts module within the interactive multimodal alignment module.
0.12 (c) Center-specific MoME Training
We adopt the sparse MoE training mechanisms[18] as the backbone structure of our proposed center-specific, multicenter training framework. The proposed MoME module consists of multiple center-specific router network $G^{c}$ and expert modules. During training, we utilized center flag $c$, where $c\in\{A,B,C\}$ for representing each Center A, B, and C. Given each data along with this center flag $c$, the corresponding center-specific router network $G^{c}$ is selectively activated during both training and inference, and it selects the top-$k$ expert modules $E$, and outputs of each expert module are then weighted by the output of each center-specific router network.
@@
0.13 (e) Center-specific MoME Inference
@@
0.14 (d) Closed center MoME fine-tuning
@@
0.15 Details of Network Training.
@@
For data pre-processing, all chest CT images, prostate labels and PTV labels were resampled to a uniform voxel spacing of 1.0 × 1.0 × 3.0 mm3. Image intensities were truncated between -200 and 250 Hounsfield units (HU) and linearly normalized to a range between 0 and 1.0. During network training, 3D patches of 384 × 384 × 128 pixels were randomly cropped to include the entire pelvic region, along with the corresponding clinical data, using a batch size of 2. For evaluation, the full 3D CT volumes were processed with a sliding window approach, using the same patch size of 384 × 384 × 128 pixels. Throughout training, the entire large language model (LLM) was frozen, while the image encoder/decoder modules, interactive alignment modules, and text prompts were set as trainable parameters. The loss function combined binary cross-entropy and Dice loss, with equal weights of 1.0. The network was optimized using the AdamW optimizer [35], with an initial learning rate of 0.0001, for 100 training epochs. For fine-tuning the network, we optimized only the image decoder, interactive multimodal alignment modules and text prompts, while keeping the image encoder module and LLM frozen. The learning rate was reduced to 0.00001, and the network parameters were optimized for up to 500 fine-tuning epochs. The network was implemented using the open-source library MONAI [^2] https://monai.io/. All experiments were conducted using PyTorch [36] in Python, leveraging CUDA 11.4 on a single NVIDIA RTX A6000 48GB GPU. For in-house model fine-tuning, we further utilized a single NVIDIA A100 40GB GPU.
For multicenter training, we utilized the entire Center A training dataset, combined with 1-shot, 2-shot, and 3-shot samples for each trial from Centers B and C with 1-shot validation samples. The few-shot samples were randomly selected based on Prostate-Specific Antigen (PSA) clusters, where the clusters were categorized as follows: 0 – PSA values below 5.0; 1 – PSA values below 10.0; 2 – PSA values below 20.0; 3 – PSA values below 30.0; and 4 – PSA values above 30.0. For fine-tuning the pre-trained multimodal MoME for being adapted to each external center, we utilized 1-shot, 2-shot, and 3-shot samples from each center D or E within their server with 1-shot validation samples.
0.16 Details of Evaluation.
@@
To quantitatively assess PTV delineation performance, we calculated the Dice coefficient (Dice) and Intersection over Union (IoU) for each patient’s PTV delineation result. We further calculated the 95th percentile of the Hausdorff Distance (95-HD) [37] to evaluate spatial discrepancies between the ground-truth and predicted contours. For reporting 95-HD, all measured distances in pixel units were adjusted according to the original pixel resolution and reported in centimeters (cm). To further evaluate center-specific PTV delineation strategies, we defined the gross tumor volume (GTV)-to-planning target volume (PTV) ratio (GPR) as the total volume of GTV divided by the total volume of PTV volume.
0.17 Statistics & Reproducibility.
@@ Acknowledgement
J.S.K. is a shareholder and employee of Oncosoft Inc, which may benefit from the research results presented in this paper. This potential conflict of interest has been disclosed and managed according to institutional policies.
@@
Supplementary Figure 2: Comparison of GTV-to-PTV ratio (GPR) across risk groups for each center by model. The vision-only model consistently demonstrates lower GPR values across all risk groups, showing significant deviation from the ground truth distribution and failing to capture the clinical context. In Centers A and B, the multi-modal and MoME models exhibit similar trends for both (a, c) low/intermediate-risk and (b, d) high/very high-risk groups. However, in Center C, where differences in data acquisition settings, patient distribution, and clinical practice patterns are pronounced, the MoME model more closely aligns with the ground truth GPR distribution compared to the multi-modal model, a trend observed in both (e) low/intermediate-risk and (f) high/very high-risk groups.
@@
Supplementary Figure 3: Comparison of GTV-to-PTV ratio (GPR) across risk groups when using each center-specific router. For Center A, both (a) low/intermediate-risk and (b) high/very high-risk groups show the closest alignment with the ground truth GPR distribution when using the Center A expert router, while the Center B expert router produces similar results. In contrast, using the Center C expert router leads to the largest deviation, effectively highlighting the similarities and differences in practice patterns across institutions. For Center B, the (c) low/intermediate-risk and (d) high/very high-risk groups also show consistent alignment with Center B’s original practice when using the Center B expert router, with a similarly close match from the Center A expert router, whereas the Center C expert router again leads to a significant increase in GPR. In Center C, both (e) low/intermediate-risk and (f) high/very high-risk groups display a distinct pattern, with higher GPR values that reflect the center’s less frequent use of PNI and tighter PTV margins.
@@
Supplementary Figure 4: Comparison of GTV-to-PTV ratio (GPR) across risk groups for the closed center D. In both (a) low/intermediate-risk and (b) high/very high-risk groups, the MoME model demonstrates a closer alignment with the ground truth distribution compared to the multi-modal as well as the vision only models. This trend becomes more pronounced as the number of examples increases with 1-shot, 2-shot, and 3-shot learning. Notably, in the high-risk group, the GPR distribution produced by the MoME model nearly matches the ground truth with just three-shot fine-tuning.
@@
Supplementary Figure 5: Comparison of GTV-to-PTV ratio (GPR) across risk groups for the closed center E. (a) low/intermediate-risk and (b) high/very high-risk groups, the MoME model demonstrates the most similar distribution with the ground truth distribution compared to the multi-modal as well as the vision only models. The distribution gets more similar to the ground truth as the few-shot tuning samples get increased to 3-shot learning, specifically in the low & intermediate group.

### v2

@@ 0.5 Ablation studies in MoME training strategy
The superiority of our multimodal MoME training is demonstrated through both quantitative analyses using traditional metrics like the Dice coefficient and qualitative GPR analysis, which converts center-specific practice patterns into comparable distributions. This analysis showed that the MoME framework effectively aligns with individual center practices. Fig. 3(a) reinforces these findings, showing that center-specific routers enable the model to closely adapt to each center’s treatment patterns, with similar centers exhibiting comparable trends. Our method proves highly adaptable in clinical settings with restricted data sharing but necessary adaptation to new data distributions. As shown in Table 4, MoME enables versatile zero-shot inference scenarios, allowing selection of the most relevant center-specific router using only a few sample test datasets. This addresses long-standing challenges in deploying AI models clinically, demonstrating that the most relevant center-specific inference significantly outperforms the least relevant one and surpasses traditional vision-only and vanilla multimodal inference. Furthermore, one- to three-shot fine-tuning using the selected center-specific router network with MoME outperforms traditional AI training mechanisms for optimizing pre-trained models in deployed clinical settings (Fig. 4). This approach is particularly valuable for real-world applications with limited sample datasets. Experimental results with varying dataset sizes suggest that our MoME fine-tuning approach efficiently adapts to each institution’s treatment strategies with minimal multicenter data, especially for complex tasks like radiotherapy target volume delineation. This demonstrates the potential for AI-based models to be effectively deployed in clinical scenarios, offering a promising solution for adapting to diverse patient populations and treatment protocols efficiently.
Despite advantages over traditional AI methods, our multicenter training mechanism has limitations. The primary drawback is reliance on limited samples for fine-tuning, potentially restricting model robustness in complex clinical environments. Future research should incorporate larger, more diverse datasets to enhance generalizability. Our method showed relative effectiveness in closed-center fine-tuning, but overall Dice metric performance remained below 0.7, with only marginal improvement compared to multimodal AI. The closed-center setting diverges from MoME’s core rationale of cross-center dataset integration. However, assessment of center A-C synergistic training revealed notable performance gains with joint dataset use. This suggests utilizing publicly accessible datasets could significantly improve closed-center fine-tuning, warranting further exploration. Additionally, evaluation metrics remain constrained. The GPR metric, while useful for visualizing practice pattern differences, is an indirect measure capturing only certain aspects. For instance, in prostate cancer target delineation, the inclusion of seminal vesicles may not be fully reflected in GPR analysis. To address this, future work will incorporate multiple metrics into the MoME model for comprehensive fine-tuning performance evaluation.
Lastly, a crucial next step is addressing inter-clinician variability, expanding beyond institutional differences. This approach aims to account for and mitigate variations in individual clinicians’ strategies and practices, potentially enhancing the model’s adaptability and clinical usability across diverse settings. Implementation will involve an interactive correction system, refining AI-driven outputs through minimal parameter adjustments based on center-specific learning paths. This approach aligns the model with diverse clinical perspectives and supports continual learning reflecting evolving treatment philosophies. Incorporating clinician-specific adaptability could lead to more personalized patient care and improve the model’s generalizability in diverse real-world applications.
@@
In conclusion, our study marks a significant step toward enabling collaboration on multicenter datasets despite challenges associated with large-scale data collection and practical constraints across institutions. The proposd MoME offers an effective method for addressing variability in radiotherapy target delineation practices. Our approach significantly outperformed recent AI training methods, demonstrating strong generalization to diverse clinical settings and adaptability to distribution shifts. This adaptability further positions the MoME framework as a promising candidate for multicenter collaborations, especially in addressing complex and often debated clinical decision-making tasks by fostering collaborative synergy and aligning with unique institutional strategies.
@@ Methods
0.6 Ethic committee approval
The hospital data deliberately collected for this study were ethically approved by the Institutional Review Boards (IRB) of the Department of Radiation Oncology at Yonsei Cancer Center, Department of Radiation Oncology at Yongin Severance Hospital, and Department of Radiation Oncology at Gangnam Severance Hospital (IRB numbers 4-2023-0179, 9-2023-0161, and 3-2023-0396, respectively), Department of Radiation Oncology at Mayo Clinic (IRB number 13-005709), and Massachusetts General Hospital (MGH) (IRB number 2022P001512). The requirement for informed consent was waived due to the retrospective nature of the study.
0.7 Target volume delineation in prostate cancer radiotherapy
In radiation oncology, treatment target volumes are classified as Gross Tumor Volume (GTV), Clinical Target Volume (CTV), and Planning Target Volume (PTV). GTV represents the visible tumor, aligning with the goal of traditional segmentation to delineate observable image regions. CTV, while sometimes derived directly from GTV when a gross tumor is present, typically encompasses areas at risk for microscopic disease. Defining CTV requires consideration of multiple clinical factors, including tumor type, histological findings, cancer stage according to the Tumor, Node, Metastasis (TNM) classification system, and, in specific cases, patient age and performance status. PTV further extends CTV to account for uncertainties in patient setup and positioning.
@@ 0.7 Target volume delineation in prostate cancer radiotherapy
In definitive radiotherapy, the target volume typically includes the prostate, seminal vesicles, and areas of suspected extracapsular extension [31]. For postoperative radiotherapy, the target encompasses the prostate bed and the seminal vesicle bed, with consideration of the surrounding anatomical structures based on the surgical field [32]. If lymph node involvement is confirmed, or if the patient falls within the high-risk or very high-risk groups according to NCCN guidelines [22], pelvic nodal irradiation (PNI) is recommended, even in the absence of radiographic evidence of lymph node metastasis. Despite the existence of general principles guiding target volume delineation, there is considerable variability in practice, particularly in the specifics of margin sizes and the inclusion of adjacent structures in cases of suspected locoregional invasion. Intermediate-risk patients with unfavorable factors may also receive PNI depending on institutional protocols, although this is not universally applied.
To reduce variability in target volume delineation, multiple consensus guidelines have been developed for both definitive and postoperative radiotherapy in prostate cancer [31, 22]. However, significant inter-institutional variability persists, influenced by multifactorial elements including clinician-specific preferences, established institutional protocols, diverse educational paradigms, and region-specific practices that reflect local population demographics and disease characteristics. Complete standardization remains challenging, emphasizing the need for strategies that can adapt to the specific practices of individual institutions while maintaining high standards of care.
0.8 Dataset characteristics and clinical context
@@
0.9 Multimodal MoME framework
The detailed schematic of our proposed network is illustrated in Supplementary Fig. 1. For the image encoder/decoder for each organ extractor and PTV delineation network, we employed the 3D Residual U-Net [24]. For the large language model (LLM), we utilized the pre-trained Llama3-8B-chat [33] model. Our proposed multimodal MoME is composed of four key steps: (i) prostate segmentation, (ii) interactive multimodal alignment, (iii) center-specific MoME training, (iv) center-specific MoME inference, and (v) closed-center MoME fine-tuning.
0.10 (i) Prostate segmentation
@@ 0.9 Multimodal MoME framework
0.11 (ii) Interactive multimodal alignment
For our multimodal AI training, we adapted a large language model (LLM) to extract clinical data. As illustrated in Supplementary Fig. 1 (b), to efficiently transfer the extensive knowledge of the LLM, we employed text prompt tuning, as previously introduced in our context-aware 3D segmentation network, LLMSeg [4]. By expanding this approach for fine-grained multimodal alignment, we embeds the entire sentence of clinical data through the frozen LLM into $L$ token-wise context embeddings ${g}\in{\mathbb{R}^{L\times D}}$, where $D$ is the LLM embedding dimension. To align these context embeddings ${g}$ with the image embeddings ${f_{l}}\in\mathbb{R}^{H_{l}W_{l}S_{l}\times Ch_{l}}$, where ${f_{l}}$ is the output of the $l$-th layer of the 3D image encoder, with $H_{l}$, $W_{l}$, and $S_{l}$ representing the height, width, and slice of the image embeddings, and $Ch_{l}$ the intermediate channel dimension, we first project ${g}$ to match the dimensions of each ${f_{l}}$ using a layer-wise linear transformation. Then, these linearly projected context embeddings ${\bar{g}}_{l}\in{\mathbb{R}^{L\times Ch_{l}}}$ are subsequently processed through self-attention and cross-attention mechanisms with the image embeddings ${f_{l}}$ within two-way transformer modules of SAM [34] structure, resulting in multimodal image embeddings ${\bar{f_{l}}}\in\mathbb{R}^{H_{l}W_{l}S_{l}\times Ch_{l}}$. In contrast to the original LLMSeg, for multicenter training, we further introduced a mixture of multicenter experts module within the interactive multimodal alignment module.
0.12 (iii) Center-specific MoME training
We adopt the sparse MoE training mechanisms[18] as the backbone structure of our proposed center-specific, multicenter training framework. The proposed MoME module consists of multiple center-specific router network $G^{c}$ and shallow multi-layer perceptron neural networks (MLP)-based expert modules. During training, we utilized center flag $c$, where $c\in\{A,B,C\}$ for representing each Center A, B, and C. Given each data along with this center flag $c$, the corresponding center-specific router network $G^{c}$ is selectively activated during both training and inference, and it selects the top-$k$ expert modules $E$, and outputs of each expert module are then weighted by the output of each center-specific router network.
@@
0.13 (iv) Center-specific MoME inference
@@
0.14 (v) Closed-center MoME fine-tuning
@@
0.15 Data pre-processing and model training
@@
For data pre-processing, all chest CT images, prostate labels, and PTV labels were resampled to a uniform voxel spacing of 1.0 × 1.0 × 3.0 mm3. Image intensities were truncated between -200 and 250 Hounsfield units (HU) and linearly normalized to a range between 0 and 1.0. During network training, 3D patches of 384 × 384 × 128 pixels were randomly cropped to include the entire pelvic region, along with the corresponding clinical data, using a batch size of 2. For evaluation, the full 3D CT volumes were processed with a sliding window approach, using the same patch size of 384 × 384 × 128 pixels. Throughout training, the entire large language model (LLM) was frozen, while the image encoder/decoder modules, interactive alignment modules, and text prompts were set as trainable parameters. The loss function combined binary cross-entropy and Dice loss, with equal weights of 1.0. The network was optimized using the AdamW optimizer [35], with an initial learning rate of 0.0001, for 100 training epochs. For fine-tuning the network, we optimized only the image decoder, interactive multimodal alignment modules, and text prompts, while keeping the image encoder module and LLM frozen. The learning rate was reduced to 0.00001, and the network parameters were optimized for up to 500 fine-tuning epochs. The network was implemented using the open-source library MONAI [^2] https://monai.io/. All experiments were conducted using PyTorch [36] in Python, leveraging CUDA 11.4 on a single NVIDIA RTX A6000 48GB GPU. For in-house model fine-tuning, we further utilized a single NVIDIA A100 40GB GPU. For multicenter training, we utilized the entire Center A training dataset, combined with 1-shot, 2-shot, and 3-shot samples for each trial from Centers B and C with 1-shot validation samples. The few-shot samples were randomly selected based on Prostate-Specific Antigen (PSA) clusters, where the clusters were categorized as follows: 0 – PSA values below 5.0; 1 – PSA values below 10.0; 2 – PSA values below 20.0; 3 – PSA values below 30.0; and 4 – PSA values above 30.0. For fine-tuning the pre-trained multimodal MoME to adapt to each external center, we utilized 1-shot, 2-shot, and 3-shot samples from each center D or E within their server with 1-shot validation samples.
0.16 Evaluation
@@
To quantitatively assess PTV delineation performance, we calculated the Dice coefficient (Dice) and Intersection over Union (IoU) for each patient’s PTV delineation result. We further calculated the 95th percentile of the Hausdorff Distance (95-HD) [37] to evaluate spatial discrepancies between the ground-truth and predicted contours. For reporting 95-HD, all measured distances in pixel units were adjusted according to the original pixel resolution and reported in centimeters (cm). To further evaluate center-specific PTV delineation strategies, we defined the gross tumor volume (GTV)-to-planning target volume (PTV) ratio (GPR) as the total volume of GTV divided by the total volume of PTV.
0.17 Statistics & reproducibility
@@ Acknowledgment
There is no competing interest.
@@
Supplementary Figure 2: Comparison of GTV-to-PTV ratio (GPR) across risk groups for each center by model. The vision-only model consistently demonstrates lower GPR values across all risk groups, showing significant deviation from the ground truth distribution and failing to capture the clinical context. In Centers A and B, the multi-modal and MoME models exhibit similar trends for both (a, c) low/intermediate-risk and (b, d) high/very high-risk groups. However, in Center C, where differences in data acquisition settings, patient distribution, and clinical practice patterns are pronounced, the MoME model more closely aligns with the ground truth GPR distribution compared to the multi-modal model, a trend observed in both (e) low/intermediate-risk and (f) high/very high-risk groups. For (a-f), the center line indicates the median, and the bounds of the box represent the interquartile range (from the lower quartile to the upper quartile).
@@
Supplementary Figure 3: Comparison of GTV-to-PTV ratio (GPR) across risk groups when using each center-specific router. For Center A, both (a) low/intermediate-risk and (b) high/very high-risk groups show the closest alignment with the ground truth GPR distribution when using the Center A expert router, while the Center B expert router produces similar results. In contrast, using the Center C expert router leads to the largest deviation, effectively highlighting the similarities and differences in practice patterns across institutions. For Center B, the (c) low/intermediate-risk and (d) high/very high-risk groups also show consistent alignment with Center B’s original practice when using the Center B expert router, with a similarly close match from the Center A expert router, whereas the Center C expert router again leads to a significant increase in GPR. In Center C, both (e) low/intermediate-risk and (f) high/very high-risk groups display a distinct pattern, with higher GPR values that reflect the center’s less frequent use of PNI and tighter PTV margins. For (a-f), the center line indicates the median, and the bounds of the box represent the interquartile range (from the lower quartile to the upper quartile).
@@
Supplementary Figure 4: Comparison of GTV-to-PTV ratio (GPR) across risk groups for the closed center D. In both (a) low/intermediate-risk and (b) high/very high-risk groups, the MoME model demonstrates a closer alignment with the ground truth distribution compared to the multi-modal as well as the vision only models. This trend becomes more pronounced as the number of examples increases with 1-shot, 2-shot, and 3-shot learning. Notably, in the high-risk group, the GPR distribution produced by the MoME model nearly matches the ground truth with just three-shot fine-tuning. For (a-b), the center line indicates the median, and the bounds of the box represent the interquartile range (from the lower quartile to the upper quartile).
@@
Supplementary Figure 5: Comparison of GTV-to-PTV ratio (GPR) across risk groups for the closed center E. (a) low/intermediate-risk and (b) high/very high-risk groups, the MoME model demonstrates the most similar distribution with the ground truth distribution compared to the multi-modal as well as the vision only models. The distribution gets more similar to the ground truth as the few-shot tuning samples get increased to 3-shot learning, specifically in the low & intermediate group. For (a-b), the center line indicates the median, and the bounds of the box represent the interquartile range (from the lower quartile to the upper quartile).